how do i solve and graph y=x^2+6x-7? i know i have to use x=-b/2a but i'm confused on exactly how this works. i've been sick and have missed 4 days of school and need help. thanks!
Answer by woolybear(13)

y=x^(2)+6x-7
To find the x-intercept, substitute in 0 for y and solve for x.
(0)=x^(2)+6x-7
Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation.
x^(2)+6x-7=(0)
Remove the parentheses around the expression 0.
x^(2)+6x-7=0
In this problem 7*-1=-7 and 7-1=6, so insert 7 as the right hand term of one factor and -1 as the right-hand term of the other factor.
(x+7)(x-1)=0
Set each of the factors of the left-hand side of the equation equal to 0.
x+7=0_x-1=0
Since 7 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 7 from both sides.
x=-7_x-1=0
Set each of the factors of the left-hand side of the equation equal to 0.
x=-7_x-1=0
Since -1 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 1 to both sides.
x=-7_x=1
The complete solution is the set of the individual solutions.
x=-7,1
To find the y-intercept, substitute in 0 for x and solve for y.
y=(0)^(2)+6(0)-7
Expand the exponent (2) to the expression.
y=(0^(2))+6(0)-7
Squaring a number is the same as multiplying the number by itself (0*0). In this case, 0 squared is 0.
y=(0)+6(0)-7
Multiply 6 by each term inside the parentheses.
y=0+0-7
Solve the equation.
y=-7
These are the x and y intercepts of the equation y=x^(2)+6x-7.
x=-7,1, y=-7